# Benchmarking Cross-Domain Audio-Visual Deception Detection

### v2

@@ 1 Introduction
Figure 1: Typical samples from different publicly deception detection datasets: Real Life Trials [1], Bag of Lies [2], MU3D [3], and Box of Lies [4]. The samples in each row are from different datasets while those in each column are with different modalities (visual vs. audio) and ground truth labels (i.e., truthful vs. deceptive). It can be seen that serious domain shifts (e.g., resolution/illumination/pose in visual faces and pitch/loudness/noise in audio) occur among these datasets.
@@
Recently, the development of automated deception detection systems using AI and machine learning techniques has gained significant attention as the existing methods have limitations in terms of reliability, accuracy, and scalability. Various multimodal datasets have been introduced, including real-life trials from court scenes [1], lab-based setups [2, 3], and game show scenarios [4]. These datasets provide a wide variety of deceptive samples from different domains, enabling researchers to examine the effectiveness of AI models on deception detection. Based on these datasets, progress has been made in deception detection techniques within specific domains [7, 6, 17]. Recent studies have utilized rich visual and audio features [18, 19, 20], such as Mel Spectrogram, emotional states, and facial action units, to enhance the performance of deception detection tasks.
@@
However, there remains a substantial research gap that needs to be addressed. Specifically, fewer studies have explored the cross-domain issue, despite the presence of significant domain shifts in public deception detection datasets. As shown in Figure 1, domain shifts are observed in both audio and visual modalities from publicly available datasets. The generalizability of the models is critical for practical applications. Therefore, such domain shifts need to be investigated in order to develop deception detection models that can be generalized across different contexts. Additionally, effective methods must be proposed to alleviate the domain shift issue by fusing both audio and visual features in a meaningful way. Addressing these issues can benefit automated deception detection systems in improving generalizability in real-world applications.
@@
To address the issue of cross-domain deception detection, we introduce a new benchmark that evaluates the generalization capacity of AI models using audio and visual features over publicly available datasets. Our benchmarking approach utilizes widely adopted audio and visual features, and we compare the single-to-single domain performance and multi-to-single domain generalization using different architectures. Specifically, for the multi-to-single setting, three domain sampling strategies, i.e., domain simultaneous, domain alternating, and domain-by-domain, are implemented to conduct cross-domain testing. To further enhance performance, we propose Multimodal Inter-Domain Gradient Matching (MM-IDGM) to improve multi-to-single domain generalization performance and an Attention-Mixer fusion method based on MLP-Mixer [21]. In our experiments, we compared single-to-single and multi-to-single domain generalization, incorporating three sampling strategies, one domain generalization method, and six fusion methods across two modalities. The benchmarking results among different features reveal that the best average fusion accuracy was 58.58% for single-to-single cross-domain generalization and 59.13% for multi-to-single cross-domain generalization. Moreover, for MM-IDGM, we achieved the highest average accuracy of 59.69% based on domain-simultaneous sampling. This benchmarking framework serves as an important tool for evaluating the effectiveness of audio-visual deception detection models in diverse contexts, which will help improve the capabilities of automated deception detection systems in real-world settings. Additionally, we hope our work will inspire further research on multimodal models that address domain shift issues. In summary, our main contributions include:
@@
  • Introducing a new benchmark for cross-domain audio-visual deception detection: We present a comprehensive benchmark that evaluates the generalization capacity of AI models using audio and visual modalities across multiple domains in the deception detection task.
  • Comparing single-to-single and multi-to-single domain generalization across different architectures: We conduct extensive experiments comparing the performance of AI models in single-to-single and multi-to-single domain generalization over different architectures and modalities inputs.
@@
  • Testing three domain sampling strategies for multi-to-single domain generalization: We propose and evaluate three distinct domain sampling strategies, i.e., domain simultaneous, domain alternating, and domain-by-domain, to assess their effectiveness and offer diverse approaches to handling domain discrepancies.
@@
  • Introducing Multimodal Inter-Domain Gradient Matching (MM-IDGM) for domain generalization: We propose MM-IDGM to improve multi-to-single domain generalization by aligning gradient updates between modalities, enhancing the model’s ability to transfer knowledge across different domains.
@@
  • Proposing the Attention-Mixer fusion method for enhanced multimodal performance: We introduce the Attention-Mixer fusion method that combines MLP-Mixer and self-attention layers to capture both intra- and inter-modal interactions better, leading to improved fusion performances in audio-visual deception detection tasks.
@@ 2.1 Cues to Deception
Figure 2: Main network and method. (a) Model architecture. Visual modality includes face and behavior inputs. Audio modality includes Mel Spectrogram input. The features obtained by the respective encoders. The fusion methods include score fusion and feature fusion. (b) Domain sampling strategies. Domain-simultaneous: each batch consists of samples from multiple sources. Domain-alternating: each batch is alternatively sampled from multiple sources. Domain-by-domain: the batches are sampled from one source and then from another.
@@ 3.1 Audio and Visual Feature Learning
As shown in Fig. 2, given a detected RGB face image as input $X_{f}$, the deep features $F_{f}$ could be extracted via face encoder networks $\mathcal{E}_{f}$ (e.g., ResNet18 [35]). Similarly, behavior inputs such as the AU and/or affect features $X_{b}$ are encoded by OpenFace [36] or affect model (e.g., EmotionNet [37]) $\mathcal{E}_{b}$ to output behavior features $F_{b}$. Note that we regard both face frames and behavior features as the visual modality but differentiate them in this work as they have different types of information and representations. Given audio input $X_{a}$ (either Mel Spectrogram [38] or waveforms), audio features $F_{a}$ are extracted through audio encoder $\mathcal{E}_{a}$. The corresponding classification heads for face frames ($\mathcal{H}_{f}$), behavior features ($\mathcal{H}_{b}$), and audio features ($\mathcal{H}_{a}$) output the prediction logits $\hat{Y}_{f}$, $\hat{Y}_{b}$, and $\hat{Y}_{a}$, respectively. The fusion head $\mathcal{G}$ takes $F_{f}$, $F_{b}$, and $F_{a}$ as input. $\mathcal{G}$ is determined by the actual fusion method, e.g., liner layer, transformer layers, MLP, etc. The output logit of $\mathcal{G}$ is denoted by $\hat{Y}_{g}$. Therefore, the audio and visual learning process can be denoted as follows:
@@ 4.1 Databases and Metrics
Datasets. We benchmarked the cross-domain generalization performance based on four publicly available datasets. Real Life Trials [1] dataset is a popular real-world dataset collected from public court trials, which consists of 121 videos including 61 deceptive and 60 truthful video clips. As it is a real-world dataset, the Real Life trial dataset has more noise on both the video and audio. We filtered out some corrupted videos and obtained 108 videos (54 truthful and 54 deceptive) with 58 subjects for our experiments. Bag of Lies [2] is a multimodal dataset collected from well-controlled lab-based scenarios, where video, audio, EEG, and gaze data are collected. It has 35 subjects, 163 truthful and 162 deceptive video clips. The backgrounds for the videos are relatively clean and it is less noisy. MU3D [3] has 320 video clips and 80 subjects that cover different races and genders. It is also a lab-based dataset that uses the personal description paradigm to stimuli real-world cases. Each participant tells a positive truth, a positive lie, a negative truth, and a negative lie. Box of Lies [4] is a deception dataset collected from an online gameshow, which has 25 videos and 26 participants (6 male and 20 female). The full video set contains 29 truthful and 36 deceptive rounds of games. However, the quality of the original Box of Lies dataset is not satisfactory. The visual (the face of the participant) and audio from many clips are not matching due to the frequent changes of viewpoints. To perform a fair comparison, we preprocessed and cleaned the Box of Lies dataset. After preprocessing, 101 video clips were extracted for testing. Some of the typical samples from these datasets are shown in Fig. 1.
@@
Evaluation Metrics. In this work, we followed the widely adopted metric, binary classification accuracy (%), for experimental evaluation. The deceptive clips were labeled as 1 and truthful clips were labeled as 0.

### v3

@@ 1 Introduction
Figure 1: Some samples from six publicly available deception detection datasets: Real-Life Trial [1], Bag of Lies [2], MU3D [3], Box of Lies [4], DOLOS [5], and MDPE [6]. Each row shows one dataset, with columns grouped by modality (visual or audio) and ground truth (truthful vs. deceptive). The visual samples demonstrate domain-specific differences such as resolution (e.g., high in Bag of Lies, low in Real-Life Trial), illumination, and subjects’ ethnicities. The audio spectrograms reveal distinct patterns: controlled datasets (MU3D, Bag of Lies, MDPE) show sharp harmonics and silent gaps, signifying stable pitch and low noise. Real‑Life Trial and Box of Lies display blurred harmonics and speckled noise, indicating high ambient interference. DOLOS shows a clear pitch with slightly elevated background noise, reflecting less controlled conditions.
@@
Recently, the development of automated deception detection systems using AI and machine learning techniques has gained significant attention as the traditional methods mentioned above have limitations in terms of reliability, accuracy, and scalability. Various multimodal datasets have been introduced, including real-life trials from court scenes [1], lab-based setups [2, 3, 6], and game show scenarios [4, 5]. These datasets provide a wide variety of deceptive samples from different domains, enabling researchers to examine the effectiveness of AI models on deception detection. Based on these datasets, progress has been made in deception detection techniques within individual datasets where models are trained and evaluated on the same domain [9, 8, 19, 20]. Recent studies have utilized rich visual and audio features [21, 22, 23, 19, 24, 7, 25], such as Mel Spectrogram, emotional states, and facial action units, to enhance the performance of deception detection tasks.
@@
However, there remains a substantial research gap that needs to be addressed. Specifically, fewer studies have explored the cross-domain issue, despite the presence of significant domain shifts in public deception detection datasets. As illustrated in Figure 1, the publicly available datasets used in deception detection exhibit substantial domain shifts across both visual and audio modalities. The differences include video resolution, illumination, subjects’ ethnicities in visuals, signal-to-noise ratios (SNRs), pitch dynamics, and background noise in audio. We quantified these differences by summarizing statistics for visual and audio information (details in Sec. 4.1 and Table I). These domain discrepancies underline the challenge of generalizing deception detection models beyond the training domain, motivating our benchmark’s focus on cross-domain evaluation. Effective methods must be proposed to alleviate the domain shift issue by fusing both audio and visual features in a meaningful way. Addressing these issues can benefit automated deception detection systems in improving generalizability in real-world applications.
@@
To address the issue of cross-domain deception detection, we introduce a new benchmark that evaluates the generalization capacity of AI models using audio and visual features over publicly available datasets. Unlike previous works that assess models only within a single dataset, our benchmark systematically evaluates cross-domain generalization using widely adopted audio and visual features through two setups: single-to-single domain generalization and multi-to-single domain generalization. Specifically, for the multi-to-single setting, three domain sampling strategies, i.e., domain simultaneous, domain alternating, and domain-by-domain, are implemented for benchmarking. These strategies control how batch data is sampled during training from multiple source datasets, allowing models to learn either domain-invariant or domain-specific patterns. To further enhance performance, we propose Multimodal Inter-Domain Gradient Matching (MM-IDGM) to improve multi-to-single domain generalization performance and an Attention-Mixer fusion method based on MLP-Mixer [26]. In our experiments, we compared single-to-single and multi-to-single domain generalization, incorporating three sampling strategies, one domain generalization method, and six fusion methods across two modalities. The benchmarking results among different features reveal that the best average fusion accuracy was 56.82% for single-to-single cross-domain generalization and 58.88% for multi-to-single cross-domain generalization. Moreover, for MM-IDGM, we achieved the highest average accuracy of 59.02% based on domain-simultaneous sampling. This benchmarking framework fills a critical gap in the literature and serves as an important tool for evaluating the effectiveness of audio-visual deception detection models in diverse contexts, which will help improve the capabilities of automated deception detection systems in real-world settings. Additionally, we hope our work will inspire further research on multimodal models that address domain shift issues. In summary, our main contributions include:
@@
  • Introducing a new benchmark for cross-domain audio-visual deception detection: We present a comprehensive benchmark designed to evaluate the model generalization using audio and visual modalities across diverse datasets with domain shifts.
  • Comparing two evaluation protocols across different architectures: single-to-single and multi-to-single domain generalization. These two protocols enable comprehensive performance comparisons across varying levels of domain discrepancy.
@@
  • Developing and testing three domain sampling strategies for multi-to-single domain generalization: We propose and evaluate three distinct domain sampling strategies, i.e., domain simultaneous, domain alternating, and domain-by-domain, to assess their effectiveness and offer diverse approaches under cross-domain settings.
@@
  • Introducing Multimodal Inter-Domain Gradient Matching (MM-IDGM) for domain generalization: We propose MM-IDGM to improve multi-to-single domain generalization by aligning gradient updates between audio and visual modalities, enhancing the model’s ability to transfer knowledge across different domains.
@@
  • Proposing the Attention-Mixer fusion method for enhanced multimodal performance: We introduce the Attention-Mixer fusion method that combines MLP-Mixer and self-attention layers to better capture both intra- and inter-modal interactions, resulting in improved multimodal integration and cross-domain generalization.
@@ 2.1 Cues to Deception
Figure 2: Main network and method. (a) Model architecture. Visual modality includes face and behavior inputs. Audio modality includes Mel Spectrogram input. The features obtained by the respective encoders. The fusion methods include score fusion and feature fusion. (b) Domain sampling strategies. Domain-simultaneous: each batch consists of samples from multiple sources. Domain-alternating: each batch is alternatively sampled from multiple sources. Domain-by-domain: the batches are sampled from one source and then from another. $F_{f}$, $F_{b}$, and $F_{a}$ represent the face, behavior, and audio features, respectively. $L_{f}$, $L_{b}$, $L_{a}$, and $L_{g}$ denote the losses corresponding to the face, behavior, and audio classifiers, as well as the fusion layer.
@@ 3.1 Audio and Visual Feature Learning
As shown in Fig. 2, given a detected RGB face image as input $X_{f}$, the deep features $F_{f}$ could be extracted via face encoder networks $\mathcal{E}_{f}$ (e.g., ResNet18 [38]). Similarly, behavior inputs such as the AU and/or affect features $X_{b}$ are encoded by OpenFace [39] or affect model (e.g., EmotionNet [40]) $\mathcal{E}_{b}$ to output behavior features $F_{b}$. Note that we regard both face frames and behavior features as the visual modality but differentiate them in this work as they have different types of information and representations. Given audio input $X_{a}$ (either Mel Spectrogram [41] or waveforms), audio features $F_{a}$ are extracted through audio encoder $\mathcal{E}_{a}$. The corresponding classifiers for face frames ($\mathcal{H}_{f}$), behavior features ($\mathcal{H}_{b}$), and audio features ($\mathcal{H}_{a}$) output the prediction logits $\hat{Y}_{f}$, $\hat{Y}_{b}$, and $\hat{Y}_{a}$, respectively. The fusion layer $\mathcal{G}$ takes $F_{f}$, $F_{b}$, and $F_{a}$ as input. $\mathcal{G}$ is determined by the actual fusion method, e.g., liner layer, transformer layers, MLP, etc. The output logit of $\mathcal{G}$ is denoted by $\hat{Y}_{g}$. Therefore, the audio and visual learning process can be denoted as follows:
@@ 4.1 Databases and Metrics
Datasets.
TABLE I: Statistics on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), MU3D (M), DOLOS (D), and MDPE (E).
Datasets Avg Resolution Subject Ethnicity Demographics Avg SNR (dB) Std of SNR (dB)
Real-life Trial (R) $\sim$360×640 mix, unspecified 58, gender bias 29.30 5.13
Bag of Lies (B1) 720×1280 100% Indian 35, gendar bias 22.63 9.31
Box of Lies (B2) $\sim$480×640 mix, unspecified 26, (6M / 20F) 25.59 3.25
MU3D (M) 720×1280 50% Black and 50% White 80, gender balanced 37.07 4.01
DOLOS (D) 720×1280 mix, unspecified 213, gendar bias 30.72 4.97
MDPE (E) 1920×1080 100% Chinese 193, gendar bias 37.17 4.28
We benchmarked the cross-domain generalization performance based on four publicly available datasets. We selected six publicly accessible, widely used multimodal deception datasets to construct our cross-domain benchmark. Each dataset brings unique characteristics that are beneficial for evaluating generalization.
  • Real Life Trials [1] (R) is a popular real-world dataset collected from public court trials. Courtroom videos capturing genuine high-stakes deception. It consists of 121 videos including 61 deceptive and 60 truthful video clips. As it is a real-world dataset, the Real Life trial dataset has more noise on both the video and audio. We filtered out some corrupted videos and obtained 108 videos (54 truthful and 54 deceptive) with 58 subjects for our experiments.
  • Bag of Lies [2] (B1) is a multimodal dataset collected from well-controlled lab-based scenarios, where video, audio, EEG, and gaze data are collected. It has 35 subjects, 163 truthful and 162 deceptive video clips. The backgrounds for the videos are relatively clean and it is less noisy.
  • MU3D [3] (M) has 320 video clips and 80 subjects that cover different races and genders. It is also a lab-based dataset that uses the personal description paradigm to stimuli real-world cases. Each participant tells a positive truth, a positive lie, a negative truth, and a negative lie.
@@
  • Box of Lies [4] (B2) is a deception dataset collected from an online gameshow, which has 25 videos and 26 participants (6 male and 20 female). The full video set contains 29 truthful and 36 deceptive rounds of games. However, the quality of the original Box of Lies dataset is not satisfactory. The visual (the face of the participant) and audio from many clips are not matching due to the frequent changes of viewpoints. To perform a fair comparison, we preprocessed and cleaned the Box of Lies dataset. After preprocessing, 101 video clips were extracted for testing.
  • DOLOS [5] (D) is the largest game show deception detection dataset, featuring rich deceptive conversations. It consists of 1,675 clips from 213 subjects, including 899 deceptive and 776 truthful samples, with fine-grained audio-visual annotations. The dataset is notable for its higher quality compared to existing alternatives.
  • MDPE [6] (E) contains over 104 hours of deception content with 193 subjects. It supports not only deception detection but also downstream tasks such as personality recognition, emotion analysis. It is a lab-based dataset with monetary incentives to encourage naturalistic deceptive behaviors. Each video has a 5-grade truthful to deceptive labels from the subjects. This enables analysis of subject-level variation in deception behavior.
Taken together, these six datasets present diverse domains in terms of subject demographics, recording conditions, modalities, and data quality. This diversity is essential to rigorously evaluate cross-domain generalization for multimodal deception detection. Some of the typical samples from these datasets are shown in Fig. 1 and more statistics are provided in Table I.
@@
Evaluation Metrics. In this work, we followed the widely adopted metric, binary classification accuracy (%), for experimental evaluation. The deceptive clips were labeled as 1 and the truthful clips were labeled as 0.
TABLE II: The results of intra-domain testing accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), MU3D (M), DOLOS (D), and MDPE (E).
Modality & Inputs Method R to R B1 to B1 M to M D to D E to E
Visual (Gaze) MLP 59.96 53.45 52.03 54.95 51.89
Visual (AU+Gaze+Affect) MLP 62.87 55.84 55.23 57.18 55.38
Visual (Facial Landmark) LSTM 58.32 50.07 51.24 52.03 51.26
Visual (Face frames) ResNet18 70.10 57.78 55.15 65.23 56.57
Visual (Face frames) ResNet18+KNN 65.28 53.68 52.04 61.36 54.92
Visual (Face frames) ResNet18+SVM 66.52 54.21 54.36 62.17 56.08
Visual (Face frames) FFCSN 67.33 56.82 56.93 63.26 57.80
Audio (Acoustic + Prosodic) MLP 66.82 50.49 55.31 58.96 50.47
Audio (Mel spectrogram) ResNet18 75.78 55.69 56.45 60.06 52.99
Audio (Mel spectrogram) ResNet18+KNN 73.24 56.38 57.22 59.07 52.30
Audio (Mel spectrogram) ResNet18+SVM 73.25 56.88 58.04 60.11 52.69
concat 59.09 54.46 56.30 70.40 55.78
Visual (Face frames) KNN 60.22 54.83 70.21 69.50 56.22
+ Visual (AU+Gaze+Affect) SVM 61.03 55.21 69.96 70.05 55.39
Atten-Mixer (Ours) 63.05 55.45 57.59 71.55 56.18
concat 68.75 54.74 56.19 66.95 53.39
Visual (Face frames) KNN 67.33 55.32 56.62 65.98 53.16
+ Audio (Mel spectrogram) SVM 68.45 56.08 56.36 67.45 54.03
Atten-Mixer (Ours) 71.00 56.44 57.14 68.10 54.58
concat 68.88 56.25 53.27 68.97 51.39
Visual (AU+Gaze+Affect) KNN 67.92 54.23 50.32 65.23 50.46
+ Audio (Mel spectrogram) SVM 68.03 54.50 51.06 64.92 51.35
Atten-Mixer (Ours) 69.30 56.78 55.23 70.98 52.99
concat 69.33 58.43 55.24 72.70 54.98
Visual (Face frames) KNN 68.55 54.25 52.28 68.48 53.19
+ Visual (AU+Gaze+Affect) SVM 68.22 54.39 52.47 68.65 54.24
+ Audio (Mel spectrogram) Atten-Mixer(Ours) 70.52 59.41 56.88 73.85 58.17
TABLE III: The results of single-to-single cross-domain generalization accuracy (%) on benchmark datasets, Real-life Trial (R), Bag of Lies (B1), Box of Lies (B2), MU3D (M), DOLOS (D), and MDPE (E).
Modality & Inputs Method R to B1 R to B2 R to M B1 to R B1 to B2 B1 to M M to R M to B1 M to B2 D to R E to R E to B1 E to B2
Visual (AU) LSTM 48.11 - - 61.21 - - - - - - - - -
Visual (Gaze) MLP 47.22 59.98 54.15 57.23 55.42 49.06 45.38 50.35 52.22 53.08 50.16 51.45 53.07
Visual (AU+Gaze) MLP 50.77 65.35 56.87 58.88 58.42 50.94 46.73 51.69 53.47 54.79 51.29 52.33 54.68
Visual (Affect) MLP 50.46 58.42 50.31 50.47 51.49 52.19 66.36 51.08 60.40 54.58 51.36 52.07 54.88
Visual (AU+Gaze+Affect) MLP 54.46 59.41 54.37 50.47 57.43 54.69 60.75 51.69 55.45 55.41 51.89 52.77 55.45
Visual (Face frames) ResNet18 52.00 61.39 51.25 50.93 57.43 50.62 57.94 51.69 57.43 54.52 51.02 53.18 55.92
Visual (Face frames) ResNet18+GRU 53.54 63.37 52.81 57.41 59.41 51.56 46.73 52.92 55.45 54.80 52.03 55.06 56.32
Visual (Face frames) ResNet18+KNN 53.13 63.15 51.72 56.38 59.25 51.06 47.02 53.22 56.11 55.36 51.96 55.88 56.15
Visual (Face frames) ResNet18+SVM 53.66 63.28 52.13 56.04 60.24 52.15 47.88 53.26 55.67 56.36 52.14 55.39 56.17
Visual (Face frames) FFCSN 54.32 62.88 51.06 59.36 53.11 48.25 52.65 52.17 54.36 56.85 53.36 54.88 56.37
Audio (Acoustic + Prosodic) MLP 43.06 52.18 50.36 49.44 59.56 51.46 50.17 50.45 52.04 50.38 46.98 49.77 50.03
Audio (Mel spectrogram) ResNet18 46.77 53.47 52.19 50.47 66.34 50.62 54.21 51.38 55.45 53.43 50.56 52.04 53.68
Audio (Mel spectrogram) ResNet18+KNN 46.88 54.02 51.39 50.69 62.77 49.36 51.09 51.53 52.72 53.17 50.48 51.72 53.23
Audio (Mel spectrogram) ResNet18+SVM 49.98 54.02 51.36 50.45 62.17 50.33 52.04 52.48 53.66 54.25 51.26 52.33 54.05
Audio (Waveform) Wave2Vec 51.08 48.51 50.94 46.73 58.42 50.00 63.55 56.31 56.44 53.55 52.77 53.49 55.68
Modality & Inputs Method R to D R to E B1 to D B1 to E M to D M to E D to B1 D to B2 D to M D to E E to M E to D Avg
Visual (AU) LSTM 49.92 - - - - - - - - - - - 53.08
Visual (Gaze) MLP 49.98 57.46 50.21 50.55 52.02 49.87 55.36 50.11 53.29 52.41 49.32 50.21 51.98
Visual (AU+Gaze) MLP 51.25 52.33 51.64 53.33 52.47 50.66 50.71 51.37 53.68 55.22 50.04 51.42 53.21
Visual (Affect) MLP 51.32 53.35 52.56 53.47 50.98 51.42 51.34 51.08 54.37 54.56 51.21 51.49 53.25
Visual (AU+Gaze+Affect) MLP 52.38 54.24 53.19 53.98 51.32 52.26 51.20 50.25 55.08 55.02 51.26 51.49 53.84
Visual (Face frames) ResNet18 52.75 51.28 52.18 50.65 52.32 52.36 50.89 51.33 55.42 54.62 51.25 52.08 53.30
Visual (Face frames) ResNet18+GRU 52.54 52.36 53.88 50.87 52.68 53.94 51.34 52.36 55.69 55.22 52.36 51.44 53.84
Visual (Face frames) ResNet18+KNN 53.37 62.32 51.26 56.37 59.04 50.77 48.21 51.48 54.39 55.08 51.49 52.33 54.26
Visual (Face frames) ResNet18+SVM 54.15 61.28 50.36 55.49 59.68 50.28 49.47 50.28 53.66 54.88 52.74 54.29 54.44
Visual (Face frames) FFCSN 53.87 61.46 51.42 56.23 60.13 50.35 50.31 53.19 53.62 54.07 53.28 55.03 54.50
Audio (Acoustic + Prosodic) MLP 45.58 51.04 51.26 50.33 60.82 49.92 51.83 50.22 53.16 50.59 49.25 50.21 50.80
Audio (Mel spectrogram) ResNet18 50.01 52.11 50.59 50.48 60.37 51.29 53.08 53.86 54.36 53.68 50.26 51.42 52.88
Audio (Mel spectrogram) ResNet18+KNN 48.82 49.98 50.21 50.28 60.08 50.49 54.11 52.38 54.59 54.32 50.38 51.27 52.24
Audio (Mel spectrogram) ResNet18+SVM 49.05 50.31 49.89 60.49 56.75 52.98 55.32 53.98 55.84 54.28 51.26 52.38 53.24
Audio (Waveform) Wave2Vec 50.21 49.51 49.94 50.43 55.83 50.82 54.26 53.65 54.38 53.49 51.21 50.65 52.87